OpenStudy (anonymous):

Find x - 3y if x = 3 and y = -2. -9 -3 3 9

OpenStudy (anonymous):

9

OpenStudy (4n1m0s1ty):

just substitute 3 for and -2 for y: 3 - 3*(-2) = ??

OpenStudy (hoblos):

3+6=9

OpenStudy (anonymous):

if x=3 and y=2 then substitute the values in the xpression 3-3(-2)=3+6=9 the answer is 9